OpenStudy (anonymous):

please i need help Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 3x - 2y = 15 6x + 5y = 11 4x - 3y = 9 -8x + 6y = -18 5x - 3y = 16 -2x + 3y = -7

OpenStudy (anonymous):

-8x + 6y = -18 -- divide by 2 -4x + 3y = -9 now we have : -4x + 3y = -9 4x - 3y = 9 ---------------add 0 = 0 (correct) answer choice 3 has infinitely many solutions

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sure thing :)